OpenStudy (anonymous):

In two supplementary angles, the measure of one angle is 6 more than twice the measure of the other. The measures of the two angles are: (a) 28 and 62 (b) 32 and 58 (c) 58 and 122 (d) 62 and 118. Which one ?

OpenStudy (anonymous):

Other angle = x One angle = 2x + 6 Supplementary means they add to get 180 x + 2x + 6 = 180 3x + 6 = 180 3x = 174 x = 58 2x + 6 = 122

OpenStudy (anonymous):

the answer is either a b c or d